# A constant rank theorem for linear elliptic equations on the sphere with applications to the mixed Christoffel problem

### v1

@@ 1.3. The mixed Christoffel problem
Using the previous results, Theorem A and a continuity argument, we will deduce the following full rank theorem.
@@
Under the same assumptions of Theorem 1.1 (resp. Theorem 1.2) on AA and $f$, with $D={{\mathbb{S}}^{n}}$, assume moreover that $f$ verifies (8). Let $u$ be a solution of equation (13). Then
@@
As an application of Corollary 1.5, we get the following existence result for the mixed Minkowski problem.
@@
Let $\Omega_{1},\dots,\Omega_{n}$ be convex bodies in $\mathbb{R}^{n+1}$, with boundary of class $C^{4,\gamma}$, for some $\gamma>0$, and positive Gauss curvature. For every $r=1,\dots,n$, let $u_{r}$ be the support function of $\Omega_{r}$ and
@@
be the mixed co-factor matrix of $W_{1}(x),\dots,W_{n-1}(x)$. Let $f\in C^{2,\gamma}(\mathbb{S}^{n})$, with $f>0$ on $\mathbb{S}^{n}$, and assume that $f$ verifies condition (7) and condition (8). Let $\mu$ be the Borel measure on ${\mathbb{S}}^{n}$} with density $f$ with respect to the (n1)(n-1)-dimensional Hausdorff measure restricted to ${\mathbb{S}}^{n}$}.
Suppose that, for any $1\leq l\leq n-1$, and any $q=1,\ldots,n$,
@@ 1.4. The mixed Christoffel problem in the $3$-dimensional space
The rest of the paper is divided in two distinct and almost independent parts. Section 2.1 is devoted to the three-dimensional case ($n=2$), while in the final section we consider the case $n\geq 3$.

### v2

@@ 1.3. The mixed Christoffel problem
Using a continuity argument, from Theorem 1.1 it can be deduced that the solution of equation (13) is in fact a geometric solution, that is, its $1$-homogeneous extension to $\mathbb{R}^{n}$ is convex. The application of the continuity argument requires additional assumptions, which are more restrictive in the case $n\geq 3$, due to the non-linearity of condition (15) with respect to the matrix $A$. For this reason we present the result in two distinct statements.
@@
Let $n=2$. Under the same assumptions of Theorem 1.2 on AA and $f$, with $D=\mathbb{S}^{2}$, assume moreover that $f$ verifies (8). Let $u$ be a solution of equation (13). Then
$W=(u_{ij}+u\delta_{ij})>0\quad\mbox{on $\mathbb{S}^{2}$.}$
In particular, the $1$-homogeneous extension of $u$ to $\mathbb{R}^{n}$ is convex.
Corollary 1.6.
Let $n\geq 3$. Under the same assumptions of Theorem 1.1 on AA and $f$, with $D={{\mathbb{S}}^{n}}$, assume moreover that condition (15) is verified by the pair $f$ and $A_{t}$, for every $t\in[0,1]$, where
$A_{t}:=(1-t)I_{n}+tA;$
(here $I_{n}$ is the identity matrix of order $n$). Let $u$ be a solution of equation (13). Then
@@
As an application of the previous corollaries we get sufficient conditions for the existence of a solution to the mixed Minkowski problem. In view of the difference between the cases $n=2$ and $n\geq 3$, we state them separately. In particular, the result in the case $n=2$ is stated in §1.4.
@@
Let $n\geq 3$. Let $\Omega_{1},\dots,\Omega_{n}$ be convex bodies in $\mathbb{R}^{n+1}$, with boundary of class $C^{4,\gamma}$, for some $\gamma>0$, and positive Gauss curvature. For every $r=1,\dots,n$, let $u_{r}$ be the support function of $\Omega_{r}$ and
@@
be the mixed co-factor matrix of $W_{1}(x),\dots,W_{n-1}(x)$. Moreover, let
$A_{t}:=(1-t)I_{n}+tA,\quad t\in[0,1]$
Let $f\in C^{2,\gamma}(\mathbb{S}^{n})$, with $f>0$ on $\mathbb{S}^{n}$, and assume that $f$ verifies condition (7). Let $\mu$ be the Borel measure on ${\mathbb{S}}^{n}$} with density $f$ with respect to the (n1)(n-1)-dimensional Hausdorff measure restricted to ${\mathbb{S}}^{n}$}.
Suppose that, for every $1\leq l\leq n-1$, for every $t\in[0,1]$ and for every $q=1,\ldots,n$, the condition
(18) $\left(\delta_{ij}+\frac{\nabla_{i}\nabla_{j}(\frac{a_{qq}}{f})}{\frac{a_{qq}}{f}}+\frac{1}{2}\frac{\nabla_{i}a_{qq}\nabla_{j}a_{qq}}{a^{2}_{qq}}-\frac{1}{2}\sum_{\alpha,\beta=1}^{l}\frac{a^{\alpha\beta}\nabla_{i}a_{q\alpha}\nabla_{j}a_{q\beta}}{a_{qq}}\right)_{i,j=1,\dots,n}\geq 0$
is verified by the matrix $A_{t}$, for every $t\in[0,1]$.
@@ 1.4. The mixed Christoffel problem in the $3$-dimensional space
The rest of the paper is organized as follows. Section 2.1 is devoted to the proofs of all the results established in the three-dimensional case ($n=2$).
In Section 3 we consider the case $n\geq 3$.
The proof of Corollaries 1.5 and 1.6 is given in Section 4.
@@ 3.2. The differential inequality for the function $\phi$
4. Proof of Corollaries 1.5 and 1.6
As stated in the introduction, we exploit a continuity argument. Let $f$ and $A$ be as in the statement of either Corollary 1.5 or Corollary 1.6. Let
$A_{0}\equiv I_{n},\quad A_{1}=A,$
where $I_{n}$ is the identity matrix. Moreover, for $t\in(0,1)$, let
$A_{t}=(1-t)A_{0}+tA_{1}.$
Note that, for every $t\in[0,1]$, $A_{t}$ and $f_{t}$ verify the assumptions of Theorem 1.1. For $n\geq 3$ this is directly assumed in Corollary 1.6. For $n=2$ this is due to the fact that condition (16) is linear with respect to $A$, and the condition is verified by the pair $f$, AA, and by the pair $f$, $I_{n}$ (here the assumption that condition (8) holds for $f$ is used – see Remark 1.7).
For $t\in[0,1]$ let $u_{t}$ be the solution of equation (13), with $A=A_{t}$ and $f=f_{t}$, and let
$W_{t}=((u_{t})_{ij}+u_{t}\delta_{ij})_{i,j=1,\dots,n}\geq 0,\quad t\in[0,1].$
We consider the set
${\mathcal{T}}=\{t\in[0,1]\colon W_{t}>0\,\mbox{on ${{\mathbb{S}}^{n}}$}\}.$
By Theorem A, $0\in{\mathcal{T}}$, and in particular $\mathcal{T}\neq\emptyset$. Moreover, the condition $W_{t}>0$ implies that $\mathcal{T}$ is open. On the other hand, Theorem 1.1 ensures that $\mathcal{T}$ is closed. We deduce that ${\mathcal{T}}=[0,1]$, and this concludes the proof.